OpenStudy (anonymous):

To solve the system of equations below, Zach isolated x2 in the first equation and then substituted it into the second equation. What was the resulting equation?

OpenStudy (anonymous):

isolate x^2 in the first equation. what does it equal?

OpenStudy (anonymous):

x^2 = ?

OpenStudy (anonymous):

subtract the y2

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

X^2=25

OpenStudy (anonymous):

don't forget about the right side. What you do to one side of the equation you have to do to the other

OpenStudy (anonymous):

x^2=25-y^2

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

now plug in 25-y^2 for x^2 in the second equation

OpenStudy (triciaal):

@oscar55773